PartnerObserving Log for 2007 Apr 21/22

Observers: A. Grazian, R. Gredel.

Night Assistant: J. Hill

Telescope Operator: D. Gonzales

17:12 Turn on House Keeping Turn on Blue Channel

17:16 System ready. Temperature and pressure are OK.

17:18 start with 25 bias frames

lbcb.20070422.001853.fits first bias frame

lbcb.20070422.003221.fits last bias frame

17:34 Dark frame of 900s

lbcb.20070422.003426.fits

17:52 taken other 25 bias frames

lbcb.20070422.005139.fits first bias frame

lbcb.20070422.010505.fits last bias frame

18:07 Dark frame of 900s

lbcb.20070422.010732.fits

The sky is almost clear, with sparse clouds at low elevation. Probably the night is not photometric. Humidity is now below 60 percent. Dome is open, but still waiting for encoder resetting. Try to take some flats with the telescope at horizon.

19:23 LBC interface hangs. Restart with lbckill and lbcstart.

Flats are gone, probably!

19:26 Encoders are OK. Now waiting for LBC control initialization.

19:29 Preset telescope to RA=+07, DEC=+25 to take some flats. Start with SkyFlatTestG lbcb.20070422.023147.fits 6000 counts.

19:33 Start SkyFlatG5 with texp=10.0

first flat image lbcb.20070422.023345.fits

last flat image lbcb.20070422.023658.fits

19:38 Start SkyFlatG5Rot180 with texp=20.0

first flat image lbcb.20070422.023911.fits

last flat image lbcb.20070422.024301.fits

19:44 Preset to SA98 without doing image analysis. (The lookup table had already produced decent images.)

SA98_UB.ob lbcb.20070422.024735.fits till lbcb.20070422.025233.fits

19:53 SA98_Vgr.ob lbcb.20070422.025427.fits till lbcb.20070422.025740.fits

19:59 Preset to GRB070411 and do image analysis at elevation 40 deg.

20:09 grb070411.ob started lbcb.20070422.031027.fits till lbcb.20070422.032638.fits seeing=1.5arcsec

20:45 grb070411.ob lbcb.20070422.034625.fits till lbcb.20070422.040226.fits seeing=1.2arcsec

21:06 Preset to M81 and do image analysis with M81extrafoc.ob

21:17 Start M81V observations. seeing=1.1arcsec

21:31 Preset to Quasar Point1 and do image analysis

sdt_fan1_gr_1.ob lbcb.20070422.043231.fits till lbcb.20070422.044352.fits seeing=1.1arsec

21:47 sdt_fan1_U_1.ob lbcb.20070422.044710.fits till lbcb.20070422.051011.fits seeing=1.1-1.4arcsec, strong gradient in image size and ellipticity.

22:15 Image analysis on Point1 (focus correction only) and preset to Point4. Point4 has no stars bright enough for image analysis in U or R. sdt_fan1_U_4.ob Guiding failed on dither 1 (1.5 arcsec with open loop tracking). Guiding near the faint limit on dither 2 (still 1.5 arcsec). Guiding failed on dither 3 using a galaxy. Guiding OK on dither 4, although seeing is variable. lbcb.20070422.054932.fits has seeing=1.2

22:55 Image analysis with SDT_Uspecfastextra40.ob and then preset to Point5. Found that we need to add +2000 nm of Z4 to the Uspec_fastextra result in order to get the proper focus for these U-Bessel observations. This defocus also explains the gradient in image size. See Issues 506 and 540 for more explanation of these focus issues.

23:15 sdt_fan1_U_5.ob lbcb.20070422.061626.fits till lbcb.20070422.063955.fits seeing=0.6arcsec in U!!!

00:02 A1835-4.ob lbcb.20070422.070336.fits till lbcb.20070422.073144.fits do not use this for weak lensing. There are problems in the focus offset and we have changed the Z4 parameter by eye seeing=0.6-0.9 in U, but some images are elongated We want to redo it.

00:40 Repeat S1835-4.ob lbcb.20070422.074150.fits till lbcb.20070422.074727.fits

We focus with Uspec and then add +300 of Z4 at texp=120s

00:52 Preset to NGC5024. Image analysis and start UVshort. Queue page didn't have the new Ferraro OBs, so we did these, and marked the old ones done.

lbcb.20070422.081415.fits till lbcb.20070422.082709.fits

seeing=0.7-0.8

01:28 NGC5024 deep UV. till lbcb.20070422.085402.fits seeing=0.9arcsec

02:02 Preset to NGC4096. Image analysis in V band and observation in U. NGC_4096_U_photo.ob Attention is U-BESSEL. seeing=0.8

lbcb.20070422.090852.fits

02:14 NGC_4144_U_photo lbcb.20070422.091323.fits

NGC_4214_U_photo lbcb.20070422.091825.fits seeing=0.7

NGC_4605_U_photo lbcb.20070422.092514.fits

NGC_4449_U_photo lbcb.20070422.093135.fits

NGC_4490_U_photo lbcb.20070422.093603.fits seeing=0.7

NGC_4631_U_photo lbcb.20070422.094049.fits Image quality still quite good 40 minutes since the previous image analysis.

02:44 SA107_UB lbcb.20070422.094826.fits till lbcb.20070422.095328.fits

After lbcb.20070422.095043.fits the archive machine has stucked. JH is investigating the problem. Had to reset/reboot lbcarchive2=lbcdh1. Then had problems with checking /dev/hdb1. See Issue #609 for details. Lost an hour of science observing, but managed to fill some of it with photometric standards. Called Ferderico. David eventually cured it by forcing the disk check and letting it run. The images are on the archive machine as of 04:18.

SA107_Vgr lbcb.20070422.095457.fits till lbcb.20070422.095812.fits

03:01 Preset to IC4182 and do image analysis. We cannot do image analysis, since images are not transfered to the Repository.

03:19 Preset to NGC6503. Expose without image analysis. lbcb.20070422.102301.fits 1 arcsec U images without any active optics or focussing

NGC_5474_U_photo lbcb.20070422.102847.fits

NGC_5585_U_photo lbcb.20070422.103329.fits

03:44 PG1633_UB lbcb.20070422.104853.fits till lbcb.20070422.095734.fits

04:02 Problem with the oil pressure. Azimuth axis tripped off. The telescope does not accept commands. Still waiting... See Issue #610 for details. Lost about 30 min to this problem. Restarted the HBS oil and were able to continue standards and sky flats.

04:24 PG1633_Vgr lbcb.20070422.112633.fits till lbcb.20070422.112946.fits

04:30 SA110_UB lbcb.20070422.113419.fits till lbcb.20070422.113929.fits

SA110_Vgr lbcb.20070422.114112.fits till lbcb.20070422.114426.fits

04:45 Start exploring Sky Flats.

SkyFlatG_5.ob lbcb.20070422.115157.fits till lbcb.20070422.115527.fits

SkyFlatGrot180_5.ob lbcb.20070422.115825.fits till lbcb.20070422.120133.fits

SkyFlatBrot180_5.ob lbcb.20070422.120317.fits till lbcb.20070422.120606.fits

SkyFlatB_5.ob lbcb.20070422.120818.fits till lbcb.20070422.121108.fits

SkyFlatUspec_5.ob lbcb.20070422.121423.fits till lbcb.20070422.121703.fits

SkyFlatUspecrot180_5.ob lbcb.20070422.121931.fits till lbcb.20070422.122207.fits

05:22 Close the Dome. End of the night.

-- JohnHill - 21 Apr 2007 (Andrea Grazian as JH, and JH himself)
Topic revision: r10 - 21 Jul 2009, NormCushing
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback